# GaSLight: Gaussian Splats for Spatially-Varying Lighting in HDR

### v2

@@ Abstract
We present GaSLight, a method that generates spatially-varying lighting from regular images. Our method proposes using HDR Gaussian Splats as light source representation, marking the first time regular images can serve as light sources in a 3D renderer. Our two-stage process first enhances the dynamic range of images plausibly and accurately by leveraging the priors embedded in diffusion models. Next, we employ Gaussian Splats to model 3D lighting, achieving spatially variant lighting. Our approach yields state-of-the-art results on HDR estimations and their applications in illuminating virtual objects and scenes. To facilitate the benchmarking of images as light sources, we introduce a novel dataset of calibrated and unsaturated HDR to evaluate images as light sources. We assess our method using a combination of this novel dataset and an existing dataset from the literature. The code to reproduce our method will be available upon acceptance.
@@ 1 Introduction
Lighting is crucial not just for shaping an image, but also for ensuring realism and aesthetic appeal when inserting virtual content. Estimating accurate lighting representations from one or many photographs is vital for creating lifelike visuals that convincingly blend within their surroundings and convey realistic surface appearance. This is essential in numerous applications including multimedia productions, virtual reality, forensics, cultural displays, and more.
Despite its importance, current methods face significant limitations in capturing and representing lighting accurately. One reason for this is the difficulty in capturing high dynamic range (HDR), which require multiple exposures [10] or specialized apparatus [58, 41]. Most approaches [69, 36, 12] therefore attempt to convert regular, low-dynamic range (LDR) images to HDR. However, as will be demonstrated later, they primarily focus on extending the dynamic range for tonemapping and visualization purposes [62, 20, 24], rather than extrapolating the intensity of bright light sources.
Another reason is the difficulty of simultaneously estimating HDR lighting that is both spatially-varying and high frequency. Tab. 1 provides an overview of existing lighting representations that have been proposed in the literature. Existing approaches either focus on accurate energy estimation, using parametric light sources [17], which allow for near-field effects, such as positioning a discrete lamp in the scene. Alternatively, per-pixel spherical harmonics [18] or tend to struggle in modeling high-frequency lighting and cannot generate strong cast shadows. Volumetric representations [32, 57, 66, 33] do not accurately model high-frequency lighting variations, as needed for reflections on glossy surfaces and do not readily integrate within existing rendering engine. Another prevalent representation used by lighting estimation methods is the environment map [16, 56, 55, 9, 44] (commonly referred to as IBL, for image-based lighting), which models infinitely-distant lighting as a spherical image. This representation cannot yield near-field lighting effects, such as a lamp in the scene, and model lighting as an infinitely-distant sphere. These limitations reduce the versatility of current lighting estimation methods, impacting their ability to achieve realistic lighting in scenes.
@@
In this work, we introduce GaSLight, a novel method that addresses these challenges by capturing the high-frequency and spatially-varying lighting of a scene. It allows for near-field lighting effects, and readily integrates within existing rendering engines. GaSLight employs HDR Gaussian Splats as light representation, which is reconstructed from a set of LDR images. As opposed to HDRGS [6] (or HDR-NeRF [27]), our approach does not require alternating exposures to recover the HDR—rather we rely on a novel, specially-designed LDR-to-HDR network based on powerful diffusion models. Compared to previous NeRF-based HDR estimation approaches [19], our method is more accurate, does not require per-scene finetuning, and is much faster.
In summary, our contributions are threefold. First, we present a framework for using Gaussian Splats as a spatially-varying HDR light representation that can be estimated from a few images of a scene. This light representation can readily be used in a 3D renderer, enabling accurate and versatile lighting effects. Second, at the heart of our framework is a novel LDR-to-HDR estimation model, which leverages the strong priors in diffusion models to plausibly estimate the intensity of light sources. Finally, we compare our method on existing HDR datasets from the literature, and show that our methods best predicts the intensity of bright light sources. We also demonstrate its applicability for virtual object insertion where virtual objects can be inserted at multiple locations in the scene and relit appropriately in a multi-view setting.
@@ 2 Related work
More recently, generative imaging has been utilized in inverse tonemapping, with GlowGAN [60] being the first to recover a multi-exposure bracket, as in traditional HDR capture [10]. Exposure Diffusion [3] improves this idea by adding exposure constraints to a pretrained diffusion model to generate relative exposures. While it capitalizes on the robust priors found in diffusion models, this method is restricted to pixel-space diffusion. Subsequently, LeDiff [62] finetunes a latent diffusion model [50] to determine relative exposures and merge them into an image with an extended dynamic range. Although this method shows promising results with latent-space diffusion for inverse tonemapping, it remains constrained by the decoder’s restricted dynamic range, which does not fully represent the richness of natural lighting. Our method draws inspiration from recent diffusion-based HDR estimation to propose an iterative dynamic range expansion scheme, enabling a significantly greater increase in dynamic range than current approaches allow.
@@
The way 3D content is represented has evolved significantly over time, addressing different aspects such as geometry and appearance. Traditional methods often relied on textured 3D meshes [15] as the primary means to represent scene geometry and appearance. Implicit radiance functions, such as those introduced in Neural Radiance Fields (NeRFs) [38], have revolutionized 3D scene representations by leveraging neural networks to represent complex geometries and appearances, successfully modeling complex, view-dependent effects. HDR-NeRF [27], HDR-HexPlane [68], and others [28, 39, 65] have extended this framework to handle HDR content, allowing for the reconstruction of scenes with an extended dynamic range. PanoHDR-NeRF [19] combine an LDR-to-HDR network trained for 360° panoramas with an omnidirectional NeRF representation to capture the spatially-varying HDR light field of an indoor scene.
Later, explicit radiance functions based on 3D Gaussian Splatting [29] were proposed. HDR-GS [6] and CineGR [61], leveraged multi-exposure input photographs to capture an extended dynamic range, similarly to their predecessor HDR-NeRF. As opposed to these methods, we focus on the extrapolation of the bright light sources present in the scene and do so without the requirement of multi-exposure inputs. We therefore create an explicit GS-based HDR representation, allowing them to function effectively as light sources for 3D renderers.
Camera pose estimation
To capture scenes in the wild, a common first step is to estimate the relative camera pose between pictures taken of a given scene. The classical way to achieve this is through bundle adjustment [59], a method which has been revisited recently by COLMAP [52, 53]. Through time, classical feature descriptors used for matching made way to learned descriptors [11, 51], which evolved to end-to-end learned relative camera pose estimation methods such as Mickey [2], Dust3r [64] and its follow-up, Mast3r [31]. We encourage the reading of the survey provided in [63].
@@
have evolved significantly over time, each addressing different aspects of light modeling and scene illumination. Parametric lights, such as point lights or directional lights, have been foundational since the inception of computer graphics. Methods that learn to predict parametric lights from images have been proposed [17, 71, 22], providing intuitive control over lighting but often struggle with accurately representing complex environments.
Spherical harmonics offer an efficient representation for lighting in real-time applications. They have been used to model both irradiance [47, 21] and radiance [54]. Garon et al. [18] estimate spatially-varying spherical harmonics directly from a single image.
Environment maps [4] provide a non-parametric alternative for capturing complex lighting environments, as they represent an infinitely distant light source modeled as a spherical image. They have been employed in many lighting estimation methods [16, 56, 26, 55, 19, 9, 44]. Implicit volumetric lighting representations were also devised to model both indoor [57, 33] and outdoor [67] lights, or for inverse rendering purposes [34]. Their opaqueness thwarts their powerful modeling capabilities, making them challenging to interpret and edit. In contrast, using Gaussian splats as a lighting representation combines the strengths of both parametric and non-parametric approaches. By representing HDR Gaussians with light-emitting spherical harmonics, our method provides spatially-varying lighting capabilities while maintaining the flexibility and interpretability of parametric representations.
@@ 3.1 HDR light intensity estimation
Refer to caption Refer to caption
(a) training (b) inference
@@
Our HDR estimation method aims to produce a plausible hallucination of the intensities of the bright light sources present in an LDR image. As illustrated in fig. 1, our approach relies on a diffusion model $\mathcal{D}$, which accepts as input an image at a certain exposure $I_{0}$, and is trained (fig. 1a) to predict two versions of the input image at different exposures: $I_{-2}=\mathcal{D}_{-}(I_{0})$, and $I_{+2}=\mathcal{D}_{+}(I_{0})$. Here, the $I_{X}$ notation means applying a $2^{X}$ factor to the exposure of the input image $I_{0}$. The resulting exposures (fig. 1b) are then merged to a single HDR image using a conventional HDR merging algorithm [10].
@@
At inference time, the network can be applied recursively to the predicted exposures to further expand the dynamic range (i.e., $I_{+4}=\mathcal{D}_{+}(\mathcal{D}_{+}(I_{0}))$ as illustrated in fig. 1b. In practice, we repeat this process until images no longer contain (under- or over-)saturated pixels.
@@
We consider the output from our multi-exposure estimation as a sequence of exposure-bracketed images $\{I_{-2n},...,I_{+2m}\}$ [48], where $n$ and $m$ indicate the number of dark and bright exposures estimated by the network, respectively. We first linearize the images ($\gamma=2.2)$, and then apply the Debevec HDR fusion method [10] to combine the sequence into a single HDR image. We adjust the weighting function to give a weight of $1$ to pixels with intensity values in the range of $[0.2,0.8]$ in the original image $I_{0}$, ensuring unsaturated pixels are preserved as much as possible from $I_{0}$. This step prevents blurring due to the averaging of slightly different pixel values that can be generated by the diffusion model across exposures. The rest of the pixels are weighted using the original weighting function, $w\!=\!-2\left\lvert z-0.5\right\rvert\!+\!1$. This results in an image in linear, unbounded colorspace.
@@
Our method can also be naively applied to 360° panoramas by using a cube map representation and applying our method on each of the six faces independently. While this does not explicitly ensure consistency for light sources crossing the edges of the cube’s face, we observe that results obtained this way are usable and produce plausible lighting and reflections, effectively converting 360° into HDRI maps.
@@
We employ a dataset of 12,611 HDR 360° panoramas to train our light intensity estimation model, containing a mix of publicly available [45, 5] and captured data. At training time, we project the spherical panorama to an image plane, using random parameters drawn according to the distributions shown in tab. 2. Note that images are first re-exposed using [49] such that the of the log-luma channel is 0.17 prior to applying the exposure augmentation from tab. 2. We also randomly apply a tone-mapping operator between gamma, Filmic, or Hybrid-Log-Gamma to ensure the model is robust to various camera functions. We normalize the output of the tonemapping operators to [-$[\text{-}1,1]$ to fit Stable Diffusion’s original training range. We obtain the prompts from each image using LLaVA [35].
@@
We initialize our Stable Diffusion 1.5 backbone [50] with the IC-Light [73] checkpoint and fine-tune it for 415k iterations. We employ an initial learning rate of $10^{\text{-}5}$. Training took 10 days on 8 A100 GPUs with 40GB of VRAM. We perform 20 diffusion steps at inference, which takes a total of approximately 40 seconds on an Nvidia Titan V GPU.
3.2 Gaussian Splat lighting representation
@@
As illustrated in fig. 2, we first process each image in the set with our HDR extrapolation network (c.f. sec. 3.1) to obtain (linear) HDR images. Then, we apply the training process proposed by InstantSplat [14] to recover an HDR 3DGS representation of the scene (in linear space). To train the 3DGS, the following loss is used:
$\ell=\ell_{2}(\log(y^{*}+1),\log(y+1))\,,$ (1)
where $y$ and $y^{*}$ are the predicted and ground truth HDR pixel values, respectively.
@@ 4 Evaluation
In this section, we first discuss the datasets used for evaluation, followed by an evaluation of our light source intensity estimation model against previous LDR-to-HDR methods. Lastly, we demonstrate that our 3D lighting framework better represents lighting conditions in natural scenes than previous work.
@@ 4.1 Data
Figure 3: Comparison of images from the (a) Si-HDR [23] and (b) BtP-HDR [5] datasets. Image crops are shown on the top row, while the bottom row shows the corresponding plot of the luminance, showing of (a) a saturated sample from the Si-HDR dataset; and (b) an unsaturated sample from BtP-HDR. The large flat regions of maximum intenstiy in (a) correspond to saturated pixel values. the BtP-HDR dataset does not contain such saturation.
To evaluate our light source intensity estimation model from sec. 3.1, we employ the following two datasets from the literature.
@@
is comprised of 195 captures, each composed of 5 RAW exposures, which are merged into HDR stacks. They also provide a reference LDR image for each capture, which is obtained by applying a custom CRF and adding realistic camera noise modeled from the Canon 5D Mark III. The dataset depicts various indoor and outdoor scenes with light sources and saturated highlights in the frame.
Analyzing the dataset reveals that 82 captures hold saturated pixels, even at the fastest shutter speed, an example of this is shown in fig. 3. While these captures allow for expanding the dynamic range of LDR images for purposes such as re-tonemapping, the saturated captures cannot faithfully represent the dynamic range of light sources. In our evaluation, we eliminate the saturated captures, which results in 113 HDR images.
@@
The unsaturated HDR dataset proposed for evaluation in [5] was expressly conceived to evaluate the presence of bright light sources in images. Contrarily to Si-HDR, the LDR reference image of this dataset is an actual capture from a camera, instead of a simulated re-exposition from the HDR merged image. We dub this dataset BtP-HDR henceforth.

### v3

@@ Abstract
We present GaSLight, a method that generates spatially-varying lighting from regular images. Our method proposes using HDR Gaussian Splats as light source representation, marking the first time regular images can serve as light sources in a 3D renderer. Our two-stage process first enhances the dynamic range of images plausibly and accurately by leveraging the priors embedded in diffusion models. Next, we employ Gaussian Splats to model 3D lighting, achieving spatially variant lighting. Our approach yields state-of-the-art results on HDR estimations and their applications in illuminating virtual objects and scenes. To facilitate the benchmarking of images as light sources, we introduce a novel dataset of calibrated and unsaturated HDR captures to evaluate images as light sources. We assess our method using a combination of our dataset and an existing dataset from the literature. The code to reproduce our method is available at https://lvsn.github.io/gaslight/.
@@ 1 Introduction
Lighting is crucial not only for shaping an image, but also for ensuring realism and aesthetic appeal when inserting virtual content. Estimating accurate lighting representations from one or many photographs is vital for creating lifelike visuals that convincingly blend within their surroundings and convey realistic surface appearance. This is essential in numerous applications including multimedia productions, virtual reality, forensics, cultural displays, and more.
Despite its importance, current methods face significant limitations in capturing and representing lighting accurately. One reason for this is the difficulty in capturing high dynamic range (HDR), which requires multiple exposures [11] or specialized apparatus [58, 43]. As a result, most approaches [67, 36, 12] attempt to convert regular, low-dynamic range (LDR) images to HDR. However, as will be demonstrated later, they primarily focus on extending the dynamic range for tonemapping and visualization purposes [61, 20, 24], rather than extrapolating the intensity of bright light sources.
Another reason is the difficulty of simultaneously estimating HDR lighting that is both spatially-varying and high frequency. Tab. 1 provides an overview of existing lighting representations that have been proposed in the literature. Existing approaches either focus on accurate energy estimation, using parametric light sources [17], which allow for near-field effects, such as positioning a discrete lamp in the scene. Alternatively, per-pixel spherical harmonics [18] tend to struggle in modeling high-frequency lighting and cannot generate strong cast shadows. Volumetric representations [32, 57, 63, 33] do not accurately model high-frequency lighting variations, as needed for reflections on glossy surfaces, and do not readily integrate within existing rendering engines. Another prevalent representation used by lighting estimation methods is the environment map [16, 56, 55, 10, 46] (commonly referred to as IBL, for image-based lighting), which models infinitely-distant lighting as a spherical image. This representation cannot yield near-field lighting effects, such as a lamp in the scene, and model lighting as an infinitely-distant sphere. These limitations reduce the versatility of current lighting estimation methods, affecting their ability to achieve realistic lighting in scenes.
@@
In this work, we introduce GaSLight, a novel method that addresses these challenges by capturing the high-frequency and spatially-varying lighting of a scene. GaSLight employs HDR Gaussian Splats (GS) as light representation, which are reconstructed from a set of LDR images. We argue that GS are particularly well-suited to representing lighting, as they allow continuous, spatially-varying intensity representation with native support for high-frequency detail and efficient integration in renderers. As opposed to HDRGS [7] (or HDR-NeRF [27]), our approach does not require alternating exposures to recover the HDR—rather, we rely on a custom, specially-designed LDR-to-HDR network based on powerful diffusion models. Compared to previous NeRF-based HDR estimation approaches [19], our method is more accurate, does not require per-scene finetuning, and is much faster.
In summary, our contributions are threefold. First, we present a framework for using Gaussian Splats as a spatially-varying HDR light representation that can be estimated from a few images of a scene. This light representation can readily be used in a 3D renderer, enabling accurate and versatile lighting effects. Second, at the heart of our framework is a novel LDR-to-HDR estimation model, which leverages the strong priors in diffusion models to plausibly estimate the intensity of light sources. Finally, we compare our method with existing HDR datasets from the literature and demonstrate that it best predicts the intensity of bright light sources. We also demonstrate its applicability for virtual object insertion, where virtual objects can be inserted at multiple locations in the scene and relit appropriately in a multi-view setting.
@@ 2 Related work
More recently, generative imaging has been utilized in inverse tonemapping, with GlowGAN [59] being the first to recover multi-exposure bracketing, akin to traditional HDR capture [11]. Exposure Diffusion [2] builds on this by incorporating exposure constraints into a pretrained diffusion model to generate relative exposures, although it is limited to pixel-space diffusion. Subsequently, LeDiff [61] finetunes a latent diffusion model [52] to determine relative exposures and merge them into an extended dynamic range image. This method shows promising results with latent-space diffusion, but remains constrained by the decoder’s limited dynamic range, which does not fully represent the richness of natural lighting. Our method proposes an iterative dynamic range expansion scheme inspired by diffusion-based HDR estimation, enabling a significantly larger increase in dynamic range than current methods allow.
@@
Traditional methods often relied on textured 3D meshes [15] as the primary means of representing scene geometry and appearance. Implicit radiance functions, such as those introduced in Neural Radiance Fields (NeRFs) [40], have revolutionized 3D scene representations by leveraging neural networks to represent detailed geometries and appearances, successfully modeling complex, view-dependent effects. HDR-NeRF [27], HDR-HexPlane [65], and others [28, 41, 62] have extended this framework to handle HDR content, allowing for the reconstruction of scenes with an extended dynamic range. PanoHDR-NeRF [19] combines an LDR-to-HDR network trained for 360° panoramas with an omnidirectional NeRF representation to capture the spatially-varying HDR light field of an indoor scene.
Later, explicit radiance functions based on 3D Gaussian Splatting [30] were proposed. HDR-GS [7] and CineGR [60], leveraged multi-exposure input photographs to capture an extended dynamic range, similarly to their predecessor HDR-NeRF. In contrast to these methods, we focus on the extrapolation of bright light sources present in the scene, doing so without the requirement of multi-exposure inputs. We therefore create an explicit GS-based HDR representation, allowing them to function effectively as light sources for 3D renderers.
@@
have evolved significantly over time, each addressing different aspects of light modeling and scene illumination. Parametric lights, such as point lights or directional lights, have been foundational since the inception of computer graphics. Methods that learn to predict parametric lights from images have been proposed [17, 69, 22], providing intuitive control over lighting, but often struggle to accurately represent complex environments.
Spherical harmonics offer an efficient representation for lighting in real-time applications. They have been used to model both irradiance [49, 21] and radiance [54]. Garon et al. [18] estimate spatially-varying spherical harmonics directly from a single image. Similarly, spherical Gaussians are an efficient basis to produce low dimensionality environment lighting [71, 6].
Environment maps [3] provide a non-parametric alternative for capturing complex lighting environments, as they represent an infinitely distant light source modeled as a spherical image. They have been employed in many lighting estimation methods [16, 56, 26, 55, 19, 10, 46]. Implicit volumetric lighting representations were also devised to model both indoor [57, 33] and outdoor [64] lights, or for inverse rendering purposes [34] [66, 5]. Their opaqueness thwarts their powerful modeling capabilities, making them challenging to interpret and edit. Explicit spatial light representations have also been introduced  [29], but the number of sources is limited and used only to represent indirect lighting. In contrast, our method uses Gaussian splats as lighting representation, which combines spatially-varying lighting capabilities while maintaining the flexibility and interpretability of parametric representations.
@@ 3.1 HDR light intensity estimation
(a) training
(b) inference
@@
Our HDR estimation method aims to produce a plausible hallucination of the intensities of the bright light sources present in an LDR image. As illustrated in fig. 1, our approach relies on a diffusion model $\mathcal{D}$, which accepts as input an image at a certain exposure $I_{0}$, and is trained (fig. 1a) to predict two versions of the input image at different exposures: $I_{-2}=\mathcal{D}_{-}(I_{0})$, and $I_{+2}=\mathcal{D}_{+}(I_{0})$. Here, the $I_{X}$ notation means applying a $2^{X}$ factor to the exposure of the input image $I_{0}$. The resulting exposures (fig. 1b) are then merged into a single HDR image using a conventional HDR merging algorithm [11].
@@
At inference time, the network can be applied recursively to the predicted exposures to further expand the dynamic range (i.e., $I_{+4}=\mathcal{D}_{+}(\mathcal{D}_{+}(I_{0}))$ as illustrated in fig. 1b. In practice, we repeat this process until images no longer contain (under- or over-)saturated pixels, which takes on average roughly 5 iterations.
@@
We generate HDR images from exposure-bracketed sequences $\{I_{-2n},...,I_{+2m}\}$ [50], where $n$ and $m$ represent the number of dark and bright exposures estimated by the network, respectively. We first linearize the images by assuming a gamma $\gamma\!=\!2.2$ tonemapping, followed by the Debevec HDR fusion method [11] to combine the sequence into a single HDR image. To preserve unsaturated pixels from $I_{0}$ and reduce blurring caused by averaging diffusion model outputs across exposures, we assign weight $1$ to pixels with intensity $\in\![0.2,0.8]$. Remaining pixels use the standard weighting function, $w\!=\!-2\left\lvert z\!-\!0.5\right\rvert\!+\!1$, producing a linear unbounded color space.
@@
Our method can also be applied to 360° panoramas by using a cube map representation and processing each face independently. While this does not explicitly ensure consistency for light sources crossing the edges of the cube’s face, we observe that results are usable and produce plausible lighting and reflections, effectively converting 360° into HDRI maps.
@@
We employ a dataset of 12,611 HDR 360° panoramas to train our light intensity estimation model, containing a mix of publicly available [47, 4] and captured data. At training time, we project the spherical panorama to an image plane, using random parameters drawn according to the distributions shown in tab. 2. Note that images are first re-exposed using [51] such that the of the log-luma channel is 0.17 prior to applying the exposure augmentation from tab. 2. We also randomly apply a tone-mapping operator between gamma, Filmic, or Hybrid-Log-Gamma to ensure the model is robust to various camera functions. We normalize the output of the tonemapping operators to $\text{-}1,1]$ to fit Stable Diffusion’s original training range. We obtain the prompts from each image using LLaVA [35].
@@
We initialize our Stable Diffusion 1.5 backbone [52] with the IC-Light [72] checkpoint and fine-tune it for 415k iterations. We employ an initial learning rate of $10^{\text{-}5}$. Training took 10 days using 8 A100-40GB GPUs. At inference, we perform 20 diffusion steps, taking $\sim{}\!4$ seconds per relative EV on an Nvidia Titan V GPU.
Gaussian Splat lighting representation
@@
As illustrated in fig. 2, we first process each image in the set with our HDR extrapolation network (c.f. sec. 3.1) to obtain (linear) HDR images. Then, we apply the training process proposed by InstantSplat [14] to recover an HDR 3DGS representation of the scene (in linear space). We optimize the 3DGS using an $\ell_{1}$ loss, along with a SSIM loss on LDR pixels.
@@ 4 Evaluation
We first discuss the evaluation datasets, then compare our light source intensity estimation model to existing LDR-to-HDR methods. Finally, we show that our 3D lighting framework better represents natural scene lighting conditions than prior work.
@@ 4.1 Data
Figure 3: Comparison of images from the (a) Si-HDR [23] and (b) BtP-HDR [4] datasets. Image crops are shown on the top row, while the bottom row shows the corresponding plot of the luminance, showing (a) a saturated sample from the Si-HDR dataset; and (b) an unsaturated sample from BtP-HDR. The large flat regions of maximum intensity in (a) correspond to saturated pixel values. The BtP-HDR dataset does not contain such saturation.
To evaluate our light estimation model from sec. 3.1, we employ the following two datasets from the literature.
@@
is comprised of 195 captures, each composed of 5 RAW exposures merged into HDR stacks and a reference LDR image for each capture, which is obtained by applying a custom CRF and adding realistic camera noise modeled from the Canon 5D Mark III. The dataset depicts various indoor and outdoor scenes with light sources and saturated highlights in the frame.
Analyzing the dataset reveals that 82 captures contain saturated pixels, even at the fastest shutter speed, an example of this is shown in fig. 3. While these captures allow for expanding the dynamic range of LDR images for purposes such as re-tonemapping, the saturated captures cannot faithfully represent the dynamic range of light sources. In our evaluation, we eliminate the saturated captures, which results in 113 HDR images.
@@
The unsaturated HDR dataset proposed for evaluation in [4] was expressly conceived to evaluate the presence of bright light sources in images. Contrarily to Si-HDR, the LDR reference image of this dataset is an actual capture from a camera, instead of a simulated re-exposure from the HDR merged image. We refer to this dataset BtP-HDR henceforth.